### v3

@@ 4 Finding Small Complete Subgraphs Efficiently
    $\ell=13$. By [12, Thm. 2], it takes $O(m^{\beta(13)/2})$, where $\beta(13)=\omega(4,4,5)$. By [22, Table 3], we have $\omega(1,5/4,1)\leq 2.58$, which yields $\omega(4,4,5)=4\omega(1,5/4,1)\leq 4\times 2.58=10.32$. Consequently, $T(n,m,13)=O(m^{\beta(13)/2})=O(m^{5.16})$.
@@
    $\ell=14$. By [12, Thm. 2], it takes $O(m^{\beta(14)/2})$, where $\beta(14)=\omega(5,4,5)$. By [38, Table 1], we have $\omega(1,0.8,1)\leq 2.221$, which implies $\omega(4,4,5)=5\omega(1,0.8,1)\leq 5\times 2.221=11.105$, thus $T(n,m,14)=O(m^{\beta(14)/2})=O(m^{5.553})$.
@@
    $\ell=16$. By [12, Thm. 2], it takes $O(m^{\beta(16)/2})$, where $\beta(16)=\omega(5,5,6)$. By [38, Table 1], we have $\omega(1,6/5,1)\leq 2.5351$, which implies $\omega(5,5,6)=5\omega(1,6/5,1)\leq 5\times 2.5351\leq 12.676$, thus $T(n,m,16)=O(m^{\beta(16)/2})=O(m^{6.338})$.

### v4

@@ 4 Finding Small Complete Subgraphs Efficiently
    $\ell=13$. By [13, Thm. 2], it takes $O(m^{\beta(13)/2})$ time, where $\beta(13)=\omega(4,4,5)$. By [23, Table 3], we have $\omega(1,5/4,1)\leq 2.58$, which yields $\omega(4,4,5)=4\omega(1,5/4,1)\leq 4\times 2.58=10.32$. Consequently, $T(n,m,13)=O(m^{\beta(13)/2})=O(m^{5.16})$.
@@
    $\ell=14$. By [13, Thm. 2], it takes $O(m^{\beta(14)/2})$ time, where $\beta(14)=\omega(4,5,5)=\omega(5,4,5)$. By [1, Table 1], we have $\omega(1,0.8,1)\leq 2.2207$, which implies $\omega(5,4,5)=5\omega(1,0.8,1)\leq 5\times 2.2207=11.1035$, thus $T(n,m,14)=O(m^{\beta(14)/2})=O(m^{5.552})$.
@@
    $\ell=16$. By [13, Thm. 2], it takes $O(m^{\beta(16)/2})$ time, where $\beta(16)=\omega(5,5,6)$. By [39, Table 1], we have $\omega(1,6/5,1)\leq 2.5351$, which implies $\omega(5,5,6)=5\omega(1,6/5,1)\leq 5\times 2.5351\leq 12.676$, thus $T(n,m,16)=O(m^{\beta(16)/2})=O(m^{6.338})$.